The tens digit of a two-digit number is 3 more than the units digit. If the number is divided by the sum
of its digits, the quotient is 7. Find the number.

Answer by Alan3354(69443)

The tens digit of a two-digit number is 3 more than the units digit. If the number is divided by the sum
of its digits, the quotient is 7. Find the number.
--------------------------
n = 10T + U
T = U + 3
-----
(10T+U)/(T+U) = 7
(10*(U+3) + U)/((U+3 + U) = 7
(11U + 30)/(2U+3) = 7
11U+30 = 14U+21
3U = 9
U = 3
T = 6
----> 63